# Understanding and Characterizing Obfuscated Funds Transfers in Ethereum Smart Contracts

### v1

@@ A.3. Contracts Without Genuine Fund Locking
(2) Code-level Obfuscation/Backdoor Techniques and Examples
@@ 15 REQUIRE(strategiesParameters[strategyName]
Fake Process: Some functions (e.g., requestWithdraw(…) and others) appear to REQUIRE user initiation, but the key steps or conditions can be forcefully modified by the Owner. For example:
@@
Claiming to ”Protect Users”: Contracts often claim in their documentation that in the event of a security incident, functions such as emergencyWithdraw() or fulfillDeposited(…) can be activated to protect users. In the code, however, these functions are mostly restricted to onlyOwner, with no multisignature or community approval:
@@
Multiple withdrawal/transfer functions such as _withdrawFromBank(…), withdraw(…), fulfillDeposited(…), and claimTokens(…)—despite having different names, share similar logic and can all be used to extract or transfer assets.
@@
(3) Core Features Extractable from Bytecode/Tool Detection
@@
  • Abundant External CALLs and Owner Dependency: External calls such as functionDelegateCall(…) or _stableToken.safeTransfer(…) and _router.swap-ExactTokensForTokens(…) are all subject to onlyOwner control, showing that ultimate control over funds is extremely centralized.

### v2

@@ A-C S3: Contracts Without Genuine Fund Locking
Code evidence.
@@ 15 REQUIRE(strategiesParameters[strategyName]
Fake Process: Some functions (e.g., requestWithdraw(...) and others) appear to REQUIRE user initiation, but the key steps or conditions can be forcefully modified by the Owner. For example:
@@
Claiming to ”Protect Users”: Contracts often claim in their documentation that in the event of a security incident, functions such as emergencyWithdraw() or fulfillDeposited(...) can be activated to protect users. In the code, however, these functions are mostly restricted to onlyOwner, with no multisignature or community approval:
@@
Multiple withdrawal/transfer functions such as _withdrawFromBank(...) , withdraw(...) , fulfillDeposited(...) , and claimTokens(...) —despite having different names, share similar logic and can all be used to extract or transfer assets.
@@
Tool-visible signals.
@@
  • Abundant External CALLs and Owner Dependency: External calls such as functionDelegateCall(...) , _stableToken.safeTransfer(...) , and _router.swapExactTokensForTokens(...) are all subject to onlyOwner control, showing that ultimate control over funds is extremely centralized.